The Z-score uses multiple corporate income and balance sheet values to measure the financial health of a company. Assess the credit risk (z-score) of a potential borrowing firm with the following financial ratios: X1 = 0.2, X2 = 0, X3 = 0.2, X4 = 0.1, X5 = 2.0
1
Expert's answer
2020-06-17T10:26:49-0400

Details:

If,

X1=0.2 (A)

X2=0 (B)

X3=0.2 (C)

X4=0.1 (D)

X5=2.0 (E)

Z score of the potential borrowing firm can be calculated using Altmans Z score formula,

"Z-Score = 1.2A + 1.4B + 3.3C + 0.6D + 1.0E"

"Z\u2212Score\u22121.2A+1.4B+3.3C+0.6D+1.0E"

"Z-score=1.2 (0.2)+1.4(0)+3.3(0.2)+0.6(0.1)+1.0(2)"

"Z score - 0.24+0+0.06+0.66+2"

Z score = 2.96